OpenStudy (anonymous):

Write an equation in slope-intercept form of the line with an x-intercept of 6 and a y-intercept of 5.

OpenStudy (noseboy908):

slope intercept form: "y=mx+b" where m= slope, and b= y-int.

OpenStudy (anonymous):

Y=mx+5 ?

OpenStudy (noseboy908):

Reading the question again, I find that yours is unlike any question about slope-int. that I've heard before. But, my guess is y=5

OpenStudy (lgbasallote):

@noseboy908 @qwertyytrewq254 use the intercept formula (then later on change to slope intercept) \[\frac xa + \frac yb = 1\] a = x-intercept b = y-intercept

OpenStudy (anonymous):

you have to start off with a different form i'll make up an easy equation 5x + 6y = 30 if you set x = 0, you get (0,5) if you set y = 0, you get (6,0) [the parameters set forth by the question] now you must put something like this into slope-int